OpenStudy (anonymous):

x/3-7/8=2/5(x+15/4) solve for x. i am lost!!!

OpenStudy (saifoo.khan):

\[\frac{x}{3} - \frac 78 = \frac 25 \times (\frac{x+15}{4})\]

OpenStudy (saifoo.khan):

Something like that?

OpenStudy (saifoo.khan):

??

OpenStudy (anonymous):

It might by ...x+15/4

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

start with the LCM.. for the left terms.

OpenStudy (anonymous):

what??

OpenStudy (saifoo.khan):